OpenStudy (anonymous):

Which shows the prime factorization of fifty-four? 2 times 3 cubed 2 squared times 3 squared 2 cubed times 3 2 cubed times 3 squared

OpenStudy (whpalmer4):

How many times can you divide 54 by 2?

OpenStudy (anonymous):

i have no idea

OpenStudy (whpalmer4):

what is 54/2?

OpenStudy (anonymous):

26?

OpenStudy (whpalmer4):

does 2*6 = 14?

OpenStudy (anonymous):

nope

OpenStudy (whpalmer4):

okay, then 26*2 does not equal 54

OpenStudy (anonymous):

0 27?

OpenStudy (anonymous):

lol sorry im very bad at math

OpenStudy (whpalmer4):

yes, much better. can you divide 27 by 2 again, without getting a fraction or a remainder?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

o nvm no

OpenStudy (whpalmer4):

Okay, better answer :-) Can you divide 27 by 3?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

what do you get when you do so?

OpenStudy (anonymous):

3*7

OpenStudy (whpalmer4):

uh, no. 3*7 = 21

OpenStudy (anonymous):

o 3*9

OpenStudy (whpalmer4):

Okay, can you divide 9 by 3?

OpenStudy (anonymous):

sadly yes

OpenStudy (whpalmer4):

how many times in all can you divide 27 by 3 until you end up with 1? 27/3 = 9 9/3 = 3 3/3 = 1

OpenStudy (anonymous):

could you just tell me the answer lol

OpenStudy (whpalmer4):

what good would my telling you the answer do you? do you think that you just need to answer this one question, and then you'll never have to do this again?!? I have bad news for you, if so...

OpenStudy (anonymous):

eh ill find someone else to help me

OpenStudy (the_fizicx99):

 The way prime factorization works is by dividing the number as many times as you can to be left with prime numbers. Prime numbers are numbers that only can be divided by 1 and themselves.